# Towards Human-like Preference Profiling in Sequential Recommendation

### v1

@@ 6.1 Setup
We select five publicly available representative recommendation benchmark datasets for our experiments: We use five widely used real-world sequential recommendation datasets for evaluation, including MovieLens-1M Harper and Konstan (2015), Amazon-books Ni et al. (2019), Steam Kang and McAuley (2018), BeerAdvocate Leskovec and Mcauley (2012), and LastFM Celma (2010) [^3] More dataset details in Appendix E.1.
For each dataset, we apply $k$-core filtering He and McAuley (2016) to remove users and items with less than $k=5$ interactions. We construct a candidate set of 20 items from which the model selects. During training, this set is composed of 10 subsequent interactions (ensuring that the correct item is always included) and 10 randomly sampled non-interacted items. For validation and testing, the candidate set consists of the correct item plus 19 randomly sampled non-interacted items. For ML-1M, Amazon-books, and BeerAdvocate, we utilize ratings as the structured preferences to adjust the preference margins, and for Steam and LastFM, where explicit ratings are unavailable, we rely on play-hours and play-count as proxies for user structured preferences. For each user, we order the interactions chronologically, using the second-last target interaction for validation, the last one for testing, and the rest for training.
@@
We compare RecPO with two types of baseline models: (i) Traditional methods leverage sequential patterns in user behaviors to predict the next interacted item, using various modeling architectures such as recurrent neural networks (GRU4Rec Hidasi (2016)), convolutional neural networks (Caser Tang and Wang (2018)), or multi-head self-attention frameworks (SASRec Kang and McAuley (2018)). (ii) LM-based methods utilize LMs to process historical interactions and predict the next interacted item. We select two LM backbones, LLaMA3 Dubey et al. (2024) and Qwen bai2023qwentechnicalreport, and compare between the standard preference optimization baseline DPO Rafailov et al. (2024), SimPO Meng et al. (2024), a reference-free method that enhances DPO with length regularization and a fixed margin term, and S-DPO Chen et al. (2024), which adapts DPO specifically for sequential recommendation [^4] More baseline details in Appendix E.2.
Note that we exclude GPT-4o OpenAI (2024), DeepSeek-V3 et al. (2025), and other similar-sized LLMs from our baselines, as their substantially greater scale renders direct comparisons unfair and risks conflating model size with alignment quality. Our objective is to demonstrate the effectiveness of RecPO as a general preference optimization framework, which is orthogonal to model scaling and broadly applicable across backbone architectures.
@@
All experiments were performed on no more than 8 NVIDIA RTX A6000 with 48GB of VRAM. For all the preference learning approaches, we first conduct SFT for task adaptation, and then post-train models initialized from SFT checkpoints by optimizing the alignment loss in Equation 7555More implementation details in Appendix E.3.
@@
6.2 Main Results and Ablations
@@
  • SFT bridges the gap between LLMs and recommendation constraints. While LLMs possess open-world knowledge, their raw outputs often violate practical requirements (e.g., recommending non-candidate items or exceeding item limits). SFT significantly improves valid output rates, matching traditional recommenders and demonstrating the necessity to align general-purpose LLMs with specific behavioral requirements in real-world applications.
@@
  • Preference alignment further unlocks the potential of LLMs in recommendation. All preference learning methods, including our proposed RecPO, DPO, SimPO, and S-DPO, significantly outperform SFT in Hit Ratio@1, suggesting the alignment between complex preferences and ranking-centric recommendation objectives. Notably, RecPO and S-DPO surpass the standard DPO, demonstrating that multi-negative preference learning better captures nuanced user preference patterns in recommendation scenarios. Although SimPO achieves an impressive improvement in Hit Ratio@1, it exhibits a noticeable degradation in Valid Ratio compared to other approaches, which highlights the limitations of reference-free optimization in mitigating distributional discrepancies between recommendation tasks and general NLP tasks.
@@
  • RecPO enhances backbone-irrelevant preference modeling. By integrating structured preference with contextualization adaptive reward margins, RecPO universally improves Hit Ratio@1 over other LLM-based approaches across both of the language backbones. This can be attributed to its human-aligned preference modeling grounded in cognitive science principles that generalize well across instructive learning tasks. Compared to traditional recommenders on implicit feedback datasets, the performance gains of RecPO are relatively modest. We posit that this narrower gap arises from proxy-derived preference signals, which exhibit homogeneous interaction patterns that even simple traditional models can effectively capture.
Ablation Study on Margin Functions.
@@
We denote $\phi_{p}$ and $\phi_{d}$ as the scores for the preferred and dispreferred items respectively, for brevity. By default, RecPO defines the margin term $\gamma_{t}$ as the ratio of preference scores $\phi$ between positive and negative item pairs, as formalized in Equation 4. To evaluate the impact of this design choice, we introduce two alternative margin functions: (i) Log Diff, $\gamma_{r}=\lambda\log\left(\phi_{p}-\phi_{d}\right)$; (ii) Log Ratio, $\gamma_{r}=\lambda\left(\log\phi_{p}-\log\phi_{d}\right)$. As shown in Table 2, both variants outperform the strongest LLM-based recommender baseline, confirming the general utility of margin-aware optimization. RecPO ’s default ratio-based margin achieves the best overall performance by amplifying training gradients, especially when historical user ratings show low volatility. By directly contrasting $\phi_{p}$ and $\phi_{d}$ via division, it provides stronger learning signals that help the model prioritize subtle but critical preference patterns.
@@
Figure 4: Comparing between SFT, S-DPO, and RecPO from the perspectives of adhering to contextual preference (a), avoiding unfavorable items under temptation (b), identifying dis-preferred items (c), and consistently performing across varying user history lengths (d). The adherence rate and avoidance rate are defined in Section 6.3.
6.3 Probing Aligned Preference from Multiple Perspectives
We further evaluate whether the learned preferences align with human-like decision patterns by probing the model from multiple perspectives. Specifically, we ask:
  • Contextual adherence: When the candidate set includes other future highly-rated items, does the model still prioritize the correct next item, reflecting sensitivity to temporal context?
  • Aversion under temptation: When the candidate set includes future low-rated items that may appear contextually tempting, can the model avoid recommending them?
  • Dispreference recognition: When directly prompted, can the model correctly identify the item least aligned with the user’s preferences?
  • Robustness to interaction length: Does the model maintain stable performance across users with varying lengths of interaction history?
RecPO adheres to contextualized preferences.
@@ 6.3 Probing Aligned Preference from Multiple Perspectives
RecPO avoids recommending undesirable items under temptation.
@@
RecPO implicitly models aversion.
While most preference alignment focuses on promoting desirable items, an essential aspect of human-like decision-making is the ability to deliberately avoid dispreferred options. To evaluate this capacity, we construct a test set querying the model directly at inference time to identify the item least aligned with a user’s preferences, without providing any explicit supervision for aversion. This setup tests whether the model’s learned preference representation implicitly encodes negative signals alongside positive ones. As shown in Figure 4(c), RecPO consistently achieves higher aversion accuracy across both datasets, outperforming SFT and S-DPO. This suggests that RecPO internalizes a more complete structure of user preferences, capable of both attraction and avoidance. Notably, this behavior emerges without explicit aversion labels: through alignment with structured and contextualized feedback alone, RecPO learns to infer which items users are likely to reject.
RecPO performs consistently better across varying interaction lengths..
In Figure 4(d), we investigate RecPO ’s robustness to variations in historical interaction lengths using the BeerAdvocate dataset. We partition the test set into subsets based on the number of past interactions and evaluate performance within each group. RecPO exhibits sustained efficacy, consistently outperforming SFT and S-DPO with larger margins. While all models follow similar performance trends as history length increases, RecPO exhibits the greatest stability, with the lowest variance in Hit Ratio@1 (8.7% vs. 17.8% for S-DPO). These results highlight RecPO ’s adaptability to diverse context lengths—a critical trait for real-world systems where user histories vary widely.
@@ 7 Conclusion
In this paper, we address the challenge of aligning recommendation objectives with human-like decision behavior by proposing RecPO, a preference alignment framework that models structured feedback and contextual delay to emulate key aspects of human preference reasoning. RecPO introduces adaptive reward margins based on inferred preference hierarchies and interaction latency, enabling LLM-based recommenders to capture fine-grained, context-aware user preferences. Experiments across multiple benchmarks show that RecPO not only improves performance but also better reflects human-like prioritization and aversion in sequential recommendation.
More broadly, this work invites a shift in how preference is modeled: not as a fixed outcome to be predicted, but as a process to be understood. Effective recommendations should not only reflect what users have chosen, but also begin to listen: to intent, uncertainty, and change. RecPO takes a step in this direction, offering a path toward systems that move alongside users, adapting not only to what they prefer, but to how their preferences evolve.
@@ Appendix A Preliminaries
Let $\mathcal{H}_{u}=[i^{1},i^{2},...,i^{N_{u}}]$ represent the chronologically ordered sequence of historical interactions for user $u$, where each element $i^{k}$ encapsulates contextual details of the $k$-th interaction (e.g., item title, style, rating), and $N_{u}$ denotes the total number of interactions. We define $\mathcal{H}_{u}^{t}=\mathcal{H}_{u}[\colon t]$ as as the subset of interactions up to time $t$, and let $i_{p}^{t+}$ denote the next recent favorable (highly-rated) item following the interaction history at $t$. Let $\mathcal{\pi}_{\theta}$ be the LM performing the task, parameterized by $\theta$. The sequential recommendation task within the LM framework is formulated as follows: given user $u$’s interaction history $\mathcal{H}_{u}^{t}$ up to time $t$ and a candidate item set $\mathcal{C}=\{i^{(j)}\}_{j=1}^{K}$, where $\mathcal{H}_{u}^{t}\cap\mathcal{C}=\emptyset$ and $i_{p}^{t+}\in\mathcal{C}$, the model $\pi_{\theta}$ is required to predict the item that most likely be favorable to user, i.e., $i_{p}^{t+}$.
@@
.$\Bigl{.}-\beta D_{\text{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})\right]\Bigr{)},$ (8)
@@ Appendix B Derivation of Preference Distribution
$\frac{w_{i}\exp\left(-\sum_{j\in\mathcal{C}\backslash\{i\}}\gamma_{ij}\right)}{\sum_{k\in\mathcal{C}}w_{k}\exp\left(-\sum_{j\in\mathcal{C}\backslash\{k\}}\gamma_{kj}\right)}.$
@@
$\mathcal{C}_{r}=\mathcal{C}\backslash\{\sigma(1),\sigma(2),\ldots,\sigma(r-1)\}$ (14)
@@
$\frac{w_{\sigma_{(r)}}\exp\left(-\sum_{j\in\mathcal{C}_{r}\backslash\{\sigma(r)\}}\gamma_{\sigma(r)j}\right)}{\sum_{k\in\mathcal{C}_{r}}w_{\sigma_{(k)}}\exp\left(-\sum_{j\in\mathcal{C}_{r}\backslash\{k\}}\gamma_{kj}\right)}.$
@@
$\prod_{r=1}^{K-1}\frac{w_{\sigma_{(r)}}\exp\left(-\sum_{j\in\mathcal{C}_{r}\backslash\{\sigma(r)\}}\gamma_{\sigma(r)j}\right)}{\sum_{k\in\mathcal{C}_{r}}w_{\sigma_{(k)}}\exp\left(-\sum_{j\in\mathcal{C}_{r}\backslash\{k\}}\gamma_{kj}\right)}$
@@ E.2 Baselines
  • Qwen2.5 bai2023qwentechnicalreport is a recent LLM developed by Alibaba, optimized for instruction-following and multi-turn dialogue tasks.
@@ E.3 Implementation Details
All experiments were conducted on a maximum of 8 NVIDIA RTX A6000 GPUs, each with 48GB of VRAM. Our framework is implemented using Python 3.10.6, PyTorch 2.2.2, and Huggingface Transformers 4.43.3. For all LLM-based recommenders, we employ LLaMA 3.1 8B Dubey et al. (2024) and Qwen2.5-7B bai2023qwentechnicalreport as the base models for both SFT and alignment. During training, we set the learning rate to 1e-5 for all LLM-based recommenders and use the AdamW optimizer. Additionally, we apply a 5% warm-up strategy and adjust the learning rate using a cosine scheduler. A global batch size of 128 is used to balance training efficiency and memory consumption. The maximum sequence length is tailored to each dataset based on the features involved and the average title lengths. We set $\beta=1$ for all preference optimization approaches. For multi-negative preference learning, including S-DPO and our proposed RecPO, we adopt the S-DPO settings and fix the number of negatives at 3. In particular, we set the margin term in SimPO as 2 and set the parameter $\lambda$ in our method as 2. Finally, following the prompt format provided in Appendix C, we create several additional prompt templates and randomly sample one for each user sequence during training and evaluation to ensure model flexibility and generality. For all traditional recommenders, we follow the settings from previous work Chen et al. (2024) by setting the learning rate to 0.001, the batch size to 256, and using the Adam optimizer for model optimization.

### v2

@@ 5.1 Setup
We select five publicly available representative recommendation benchmark datasets for our experiments: We use five widely used real-world sequential recommendation datasets for evaluation, including MovieLens-1M Harper and Konstan (2015), Amazon-books Ni et al. (2019), Steam Kang and McAuley (2018), BeerAdvocate Leskovec and Mcauley (2012), and LastFM Celma (2010) [^3] More dataset details in Appendix E.1. For each dataset, we apply $k$-core filtering He and McAuley (2016) to remove users and items with less than $k=5$ interactions. We construct a candidate set of 20 items from which the model selects. During training, this set is composed of 10 subsequent interactions (ensuring that the correct item is always included) and 10 randomly sampled non-interacted items. For validation and testing, the candidate set consists of the correct item plus 19 randomly sampled non-interacted items. For ML-1M, Amazon-books, and BeerAdvocate, we utilize ratings as the structured preferences to adjust the preference margins, and for Steam and LastFM, where explicit ratings are unavailable, we rely on play-hours and play-count as proxies for user structured preferences. For each user, we order the interactions chronologically, using the second-last target interaction for validation, the last one for testing, and the rest for training.
@@
We compare RecPO with two types of baseline models: (i) Traditional methods leverage sequential patterns in user behaviors to predict the next interacted item, using various modeling architectures such as recurrent neural networks (GRU4Rec Hidasi (2016)), convolutional neural networks (Caser Tang and Wang (2018)), or multi-head self-attention frameworks (SASRec Kang and McAuley (2018)). (ii) LM-based methods utilize LMs to process historical interactions and predict the next interacted item. We select two LM backbones, LLaMA3 Dubey et al. (2024) and Qwen Bai et al. (2023), and compare between the standard preference optimization baseline DPO Rafailov et al. (2024), SimPO Meng et al. (2024), a reference-free method that enhances DPO with length regularization and a fixed margin term, and S-DPO Chen et al. (2024), which adapts DPO specifically for sequential recommendation [^4] More baseline details in Appendix E.2. Note that we exclude proprietary LLMs due to their lack of training access in integrating the two factors—preference intensity and temporal context—identified in our investigation. As this work is hypothesis-driven rather than solely method-focused, our objective is to validate how these factors enable effective preference modeling, which requires full control over LLMs unavailable in closed-source systems.
@@
All experiments are performed on 8 NVIDIA RTX A100 with 80GiB of VRAM. For all the preference learning approaches, we first conduct SFT for task adaptation, and then post-train models initialized from SFT checkpoints by optimizing the alignment loss in Equation 7555More implementation details in Appendix E.3.
@@
5.2 Do Preference Intensity and Temporal Context Improve Recommendation?
@@
  • SFT establishes base model capabilities. While LLMs possess open-world knowledge, their raw outputs often violate practical requirements (e.g., recommending non-candidate items or exceeding item limits). SFT significantly improves valid output rates, matching traditional recommenders and demonstrating the necessity to align general-purpose LLMs with specific behavioral requirements in real-world applications.
@@
  • Binary preference modeling provides limited gains. All preference learning methods, including our proposed RecPO, DPO, SimPO, and S-DPO, significantly outperform SFT in Hit Ratio@1, suggesting the alignment between complex preferences and ranking-centric recommendation objectives. Notably, RecPO and S-DPO surpass the standard DPO, demonstrating that multi-negative preference learning better captures nuanced user preference patterns in recommendation scenarios. Although SimPO achieves an impressive improvement in Hit Ratio@1, it exhibits a noticeable degradation in Valid Ratio compared to other approaches, which highlights the limitations of reference-free optimization in mitigating distributional discrepancies between recommendation tasks and general NLP tasks.
@@
  • Integrating preference intensity and temporal context yields substantial improvements. By integrating structured preference with contextualization adaptive reward margins, RecPO universally improves Hit Ratio@1 over other LLM-based approaches across both of the language backbones. This can be attributed to its human-aligned preference modeling grounded in cognitive science principles that generalize well across instructive learning tasks. Compared to traditional recommenders on implicit feedback datasets, the performance gains of RecPO are relatively modest. We posit that this narrower gap arises from proxy-derived preference signals, which exhibit homogeneous interaction patterns that even simple traditional models can effectively capture.
How Should Preference Intensity and Temporal Context Be Combined?
@@
We denote $\phi_{p}$ and $\phi_{d}$ as the scores for the preferred and dispreferred items respectively, for brevity. By default, RecPO defines the margin term $\gamma_{t}$ as the ratio of preference scores $\phi$ between positive and negative item pairs, as formalized in Equation 4. To investigate how these factors should be mathematically combined, we introduce two alternative margin functions: (i) Log Diff, $\gamma_{r}=\lambda\log\left(\phi_{p}-\phi_{d}\right)$; (ii) Log Ratio, $\gamma_{r}=\lambda\left(\log\phi_{p}-\log\phi_{d}\right)$. As shown in Table 2, both variants outperform the strongest LLM-based recommender baseline, confirming that incorporating these factors through any margin formulation improves performance. RecPO ’s default ratio-based margin achieves the best overall performance by amplifying training gradients, especially when historical user ratings show low volatility. By directly contrasting $\phi_{p}$ and $\phi_{d}$ via division, it provides stronger learning signals that help the model prioritize subtle but critical preference patterns.
@@
Figure 4: Comparing between SFT, S-DPO, and RecPO from the perspectives of adhering to contextual preference (a), avoiding unfavorable items under temptation (b), identifying dis-preferred items (c), and consistently performing across varying user history lengths (d). The adherence rate and avoidance rate are defined in § 5.3.
5.3 Do Models Learn Human-Aligned Preference Patterns?
Beyond quantitative performance, we investigate whether incorporating preference intensity and temporal context enables models to exhibit human-like decision patterns. We probe the learned preferences from multiple perspectives:
  • Temporal context sensitivity: When the candidate set includes other future highly-rated items, does the model still prioritize the correct next item, reflecting sensitivity to temporal context?
  • Preference intensity awareness: When the candidate set includes future low-rated items that may appear contextually tempting, can the model avoid recommending them?
  • Implicit aversion modeling: When directly prompted, can the model correctly identify the item least aligned with the user’s preferences?
  • Robustness across contexts: Does the model maintain stable performance across users with varying lengths of interaction history?
Temporal context enables immediate satisfaction prioritization.
@@ 5.3 Do Models Learn Human-Aligned Preference Patterns?
Preference intensity enables discernment under temptation.
@@
Both factors jointly enable implicit aversion modeling.
While most preference alignment focuses on promoting desirable items, an essential aspect of human-like decision-making is the ability to deliberately avoid dispreferred options. To evaluate this capacity, we construct a test set querying the model directly at inference time to identify the item least aligned with a user’s preferences, without providing any explicit supervision for aversion. This setup tests whether the model’s learned preference representation implicitly encodes negative signals alongside positive ones. As shown in Figure 4(c), RecPO consistently outperforms SFT and S-DPO with higher aversion accuracy across both datasets. This suggests that RecPO internalizes a more complete structure of user preferences, capable of both attraction and avoidance. Notably, this behavior emerges without explicit aversion labels: through alignment with structured and contextualized feedback alone, RecPO learns to infer which items users are likely to reject.
Learned patterns generalize across varying interaction contexts.
In Figure 4(d), we investigate RecPO ’s robustness to variations in historical interaction lengths using the BeerAdvocate dataset. We partition the test set into subsets based on the number of past interactions and evaluate performance within each group. RecPO exhibits sustained efficacy, consistently outperforming SFT and S-DPO with larger margins. While all models follow similar performance trends as history length increases, RecPO exhibits the greatest stability, with the lowest variance in Hit Ratio@1 (8.7% vs. 17.8% for S-DPO). These results highlight RecPO ’s adaptability to diverse context—a critical trait for real-world systems where user histories vary widely.
6 Related Work
Sequential recommendation models temporal user preferences in interaction sequences. Early methods adopt structures such as recurrent neural networks (GRU4Rec Hidasi (2016)) and self-attention mechanisms (SASRec Kang and McAuley (2018)). Recent advances integrate LLMs for their rich semantic understanding and contextual reasoning capabilities Liao et al. (2024); Bao et al. (2023); Geng et al. (2022); Yuan et al. (2023).
LLM preference alignment techniques aim to align language models’ outputs with human preferences. Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022) and DPO Rafailov et al. (2024) fine-tune models using pairwise preference data. Building on DPO, methods like IPO Azar et al. (2024), CPO Xu et al. (2024), KTO Ethayarajh et al. (2024), SimPO Meng et al. (2024), and ODPO Amini et al. (2024) further refine alignment. Most recently, S-DPO Chen et al. (2024) adapts alignment for sequential recommendation using list-wise negative items. However, these methods treat all preferences uniformly through binary comparisons, overlooking graded preference intensity and temporal context. More details are in Appendix D.
Our work differs by investigating what factors enable effective preference modeling in LLM-based recommendation, revealing that preference intensity and temporal context are critical yet overlooked dimensions.
@@ 7 Conclusion
We investigate what makes LLMs effective at modeling user preferences in sequential recommendation. Through systematic empirical study, we identify two critical factors overlooked by current binary preference modeling: preference intensity—the graded strength of user affinity or aversion—and temporal context—the immediacy of satisfaction. In contrast, prior methods that rely on binary modeling discard essential information. Motivated by these findings, we propose RecPO, which operationalizes these factors through adaptive reward margins. Our SoTA results demonstrate that preference intensity and temporal context are fundamental to effective LLM-based recommendation, with implications for human preference modeling.
@@ Appendix A Preliminaries
Let $\mathcal{H}_{u}=[i^{1},i^{2},...,i^{N_{u}}]$ represent the chronologically ordered sequence of historical interactions for user $u$, where each element $i^{k}$ encapsulates contextual details of the $k$-th interaction (e.g., item title, style, rating), and $N_{u}$ denotes the total number of interactions. We define $\mathcal{H}_{u}^{t}=\mathcal{H}_{u}[\colon t]$ as the subset of interactions up to time $t$, and let $i_{p}^{t+}$ denote the next recent favorable (highly-rated) item following the interaction history at $t$. Let $\mathcal{\pi}_{\theta}$ be the LM performing the task, parameterized by $\theta$. The sequential recommendation task within the LM framework is formulated as follows: given user $u$’s interaction history $\mathcal{H}_{u}^{t}$ up to time $t$ and a candidate item set $\mathcal{C}=\{i^{(j)}\}_{j=1}^{K}$, where $\mathcal{H}_{u}^{t}\cap\mathcal{C}=\emptyset$ and $i_{p}^{t+}\in\mathcal{C}$, the model $\pi_{\theta}$ is required to predict the item that most likely be favorable to user, i.e., $i_{p}^{t+}$.
@@
$\Bigl.-\beta D_{\text{KL}}\left[\pi_{\theta}(\mathbf{y}|\mathbf{x})\|\pi_{\text{ref}}(\mathbf{y}|\mathbf{x})\right]\Bigr),$ (8)
@@ Appendix B Derivation of Preference Distribution
[…]
@@
[…]
@@
[…]
@@ E.2 Baselines
  • Qwen2.5 Bai et al. (2023) is a recent LLM developed by Alibaba, optimized for instruction-following and multi-turn dialogue tasks.
@@ E.3 Implementation Details
All experiments were conducted on a maximum of 8 NVIDIA RTX A6000 GPUs, each with 48GB of VRAM. Our framework is implemented using Python 3.10.6, PyTorch 2.2.2, and Huggingface Transformers 4.43.3. For all LLM-based recommenders, we employ LLaMA 3.1 8B Dubey et al. (2024) and Qwen2.5-7B Bai et al. (2023) as the base models for both SFT and alignment. During training, we set the learning rate to 1e-5 for all LLM-based recommenders and use the AdamW optimizer. Additionally, we apply a 5% warm-up strategy and adjust the learning rate using a cosine scheduler. A global batch size of 128 is used to balance training efficiency and memory consumption. The maximum sequence length is tailored to each dataset based on the features involved and the average title lengths. We set $\beta=1$ for all preference optimization approaches. For multi-negative preference learning, including S-DPO and our proposed RecPO, we adopt the S-DPO settings and fix the number of negatives at 3. In particular, we set the margin term in SimPO as 2 and set the parameter $\lambda$ in our method as 2. Finally, following the prompt format provided in Appendix C, we create several additional prompt templates and randomly sample one for each user sequence during training and evaluation to ensure model flexibility and generality. For all traditional recommenders, we follow the settings from previous work Chen et al. (2024) by setting the learning rate to 0.001, the batch size to 256, and using the Adam optimizer for model optimization.